im reall yreally stuck on this one, can you please help me, so anyway, the problem is to find the value of x.
5/x = 10
HELP PLEASE!!

Answer by stanbon(75887)

5/x = 10
Multiply both sides by x to get:
5 = 10x
Divide both sides by 10 to get
x = 5/10
Simplify:
x = 1/2
===============
Cheers,
Stan H.